OpenStudy (anonymous):

Oh noes another question i only understand half of ! Find greatest common monomial factor of the terms of the polynomial. 8x3^y2^ – 12x2^y3^ + 20x2^y2^

OpenStudy (anonymous):

It just means the greatest common factor of the 3 terms

OpenStudy (anonymous):

Should be 4x^2y^2

OpenStudy (anonymous):

Ah that seems to be in one of the choices , but i must as how did u figuere it out ?

OpenStudy (anonymous):

Did you learn factorisation?

OpenStudy (anonymous):

An example 6 + 3 = 3(2 + 1) Another example 6x + 3x = 3x(2 + 1) More example 6xy + 3xy = 3xy(2 + 1) 6x^2y + 3xy = 3xy(2x + 1) 6x^2y^3 + 3x^3y^2 = 3x^2y^2(2y + x)

OpenStudy (anonymous):

thanks !